# Emergence of non-ergodic multifractal quantum states in geometrical fractals

### v3

@@ E.2 Skewness
Skewness (Sk) quantifies the asymmetry of a probability distribution [82, 83]. For the wavefunction intensity, $|\psi_{E}|_{i}^{2}$, across lattice sites, Sk reveals deviations from a symmetric spatial spread. For instance, a perfectly uniform (ideal extended) state would have $\text{Sk}=0$. In contrast, localized states highly concentrated on a few sites with most other sites having near-zero intensity would exhibit large positive skewness. Skewness therefore serves as an additional qualitative indicator of spatial inhomogeneity, complementing the quantitative details from mass exponent analysis and aiding in a deeper understanding of the eigenstate structure. It is defined as:
@@
where $|\psi_{E}|_{i}^{2}$ is the wavefunction intensity at site $i$ for an eigenstate $\Phi_{E}$, $N$ is the total number of active sites for a given system generation $g$ (i.e. $N_{g}=N$), and the brackets denotes the spatial average.
@@
Figure S16 presents the finite-size scaling (FSS) of the skewness, $\text{Sk}(N)$, for various characteristic states, obtained by plotting $\text{Sk}(N)$ against $1/\ln N$. The left panel of Fig. S16 establishes benchmark behaviors by displaying the FSS of skewness for the representative extended, localized, and multifractal states discussed in the MT. The extended ground state (GS, red circles) and the localized state (LOC, blue triangles) align with theoretical expectations: the former’s skewness values extrapolate towards $\text{Sk}_{\infty}\approx 0$ (consistent with its relatively uniform spatial spread), whereas the latter has a large, positive skewness that increases with system size $N$ (as for strong concentration on few sites). The representative non-ergodic multifractal state (NEM) also displays significant positive skewness for finite system sizes, trending towards a smaller yet distinctly positive $\text{Sk}_{\infty}$. Our analysis primarily uses this observed trend to differentiate the NEM’s behavior from both the near-zero skewness characteristic of extended states and the large, large skewness found for localized states. With this in mind, in the right panel of Fig. S16 we present the skewness FSS for the eigenstates at $E\simeq 0.000$, $E\simeq 1.000$, and $E\simeq 2.000$.
@@
Figure S16: Finite-size scaling of wavefunction intensity skewness $\text{Sk}(N)$ versus $1/\ln N$. (Left panel) Benchmark behaviors for representative extended (GS), multifractal (NEM), and localized (LOC) states. These states are extensively discussed in the Main Text. (Right panel) Skewness scaling for eigenstates at energies $E\simeq 0.000$, $E\simeq 1.000$, and $E\simeq 2.000$. Data points correspond to different system sizes and trends indicate the behavior towards the thermodynamic limit.
@@
The state at $E\simeq 0.000$ (green diamonds) is notable as it converges to a distinct, negative skewness, $\text{Sk}_{\infty}\simeq-1.055$ (determined from individual FSS fits not shown in this figure). Negative skewness typically implies that the distribution has a long tail towards lower intensity values, meaning a majority of participating sites have higher-than-average intensities, with a sharp drop-off at the very low-intensity end. The amplitude profile of this state (Fig. S17a) visually corroborates this interpretation of its asymmetry; this observation, when considered alongside the mass exponent characterization, supports its classification as weakly ergodic.
@@
The eigenstate at $E\simeq 2.000$ (cyan hexagons) is also classified as weakly ergodic. Although its skewness is relatively stable and extrapolates to a small but positive $\text{Sk}_{\infty}$, its amplitude profile (Fig. S17c) reveals a broad, non-uniform distribution of intensity across many sites, with sharp probability decays occurring only at the edges of the larger groups of voids of the SG. These features, combined with the state’s high mass exponent value, point towards its classification also as weakly ergodic.
@@
Figure S17: Intensity profiles of $|\psi|^{2}$ corresponding states at energies (from left to right): $E\simeq 0.000$, $E\simeq 1.000$, and $E\simeq 2.000$. The states for $E\simeq 0.000$ and $E\simeq 2.000$ represent case of weakly ergodic states. The state for $E\simeq 1.000$ is a case of clustered-patterned multifractality. The color intensity, indicated by the color bar, represents $|\psi|^{2}$ on a logarithmic scale.
@@
Finally, the state at $E\simeq 1.000$ (purple triangles) exhibits a clear positive extrapolated skewness. From the FSS, its $\text{Sk}(N)$ values drift towards a moderate positive value for larger system sizes. This trend, while distinct from those of the weakly ergodic states at $E\simeq 0.000$ and $E\simeq 2.000$, instead appears more similar to the benchmark NEM. Its probability distribution, while not uniform, is situated on a structured subset of sites with significant intensity fluctuations, leading to its positive skewness from extended regions of low probability. This aligns with its intensity profile (Fig. S17b), which reveals clusters of varying intensities around the $g-3$ sub-gasket blocks, supporting its characterization as a form of ’clustered-patterned multifractality’.
Appendix F Robustness against randomly displaced defects
@@
Figure S18 illustrates the evolution of the degeneracy for the special multifractal state at $E\simeq 1.000$ as a function of the number of randomly introduced defects. Contrary to the common expectation that adding defects would disrupt the fractal’s self-similar structure and lift the degeneracy of this energy level, an opposite trend is observed. The introduction of random defects can, in fact, enhance certain quantum interference patterns, leading to an increased degeneracy. The changes to its spatial intensity profile are displayed in Fig. S19, which compares representative configurations before and after the addition of these random defects.
@@
Figure S18: Change in the degeneracy $n(E)$ of the multifractal state at $E=1.000$ in a $g=6$ SG as additional random defects are introduced. The plotted values are averages over $p=10$ distinct disorder realizations.
@@
Figure S19: Intensity profiles of the multifractal state at energy $E=1.000$ on a $g=6$ SG. Left panel: The state in the pristine system, without additional random defects (ARDs). Center and right panels: Two distinct realizations of the state when 150 ARDs are introduced. While some of the clustered features may persist, the introduction of ARDs can generate new quantum interference patterns, and the intensity is often redistributed primarily towards the lower hierarchical levels of the SG structure.
Appendix G Next-nearest neighbors hopping
@@
Figure S20: The frequency correlator $\mathcal{F}(\omega,E)$ for representative non-ergodic multifractal (NEM) states, illustrating the impact of varying next-nearest-neighbor (NNN) hopping strengths $t^{\prime}$. The plots reveal how characteristics such as plateau width and slope behavior respond to changes in $t^{\prime}$. Lines serve as guides to the eye.
Figure S20 presents the behavior of $\mathcal{F}(\omega,E)$ for NEM states, specifically at energies $E\simeq 1.509$ and $E\simeq-1.287$, computed on a $g=6$ SG. The introduction of a small NNN hopping strength, up to $t^{\prime}=10^{-2}$, does not significantly modify the key qualitative features of the correlator, such as its plateau height and characteristic power-law decay exponent.
@@
Appendix H Generalization to other fractal geometries and interactions
An intriguing question arising from our findings is the extent to which the emergence of NEMs is specific to the SG, or if it represents a more general phenomenon. From the results inferred in this work, we believe NEMs can manifest in diverse fractal structures, though their characteristics would likely depend on intrinsic geometric properties such as the ramification number—a factor known to be significant in related studies on general graphs and fractals [52, 39].
In this spirit, investigating fractals embedded in three-dimensional space may offer a concrete pathway to observe NEM signatures [14]. However, performing comprehensive numerical analyses within 3D configurations presents significant computational challenges, which may currently limit extensive quantitative verification.
@@ Appendix H Generalization to other fractal geometries and interactions
H.1 Negative Vicsek fractal
Conceptually, the negative Vicsek (NV) fractal is easier to fabricate than the SG, as it requires significantly fewer defects to be placed in a controllable manner. The starting point is a uniform square lattice $\mathcal{S}$ of $N$ sites. Then, unlike in the SG case, we retain sites corresponding to voids of the fractal and remove all other sites, see Figs. S21 and  S22. We denote the set of indices of the removed sites as $\mathcal{B}\subset\mathcal{S}$.
Figure S21: Transformation of a minimal square lattice into the first-generation negative Vicsek fractal. Next-to-nearest hoppings are required to keep the lattice connected.
Figure S22: Transformation of a square lattice into the fourth-generation Vicsek saltire fractal. The final structure emerges from successive iterations of the elementary transformation, where lattice sites belonging to subset $\mathcal{B}$ are progressively replaced by defects. The resulting fractal is characterized by patching large regions of intact lattice without defects and regions with intricate defect patterns.
As in the SG case, each defect at position $r_{i}$ is modeled by a delta-peaked potential, $V_{i}=V,\delta(r-r_{i})$, with strength $V$ much larger than other energy scale in the system. However, in the NV model, we also account for next-nearest neighbor interactions, denoted by $t^{\prime}$. These additional couplings are fundamental, as they maintain connectivity between otherwise isolated regions of the fractal. The hoppings between defects and the regular sites are weak. The corresponding Hamiltonian reads:
@@ H.1 Negative Vicsek fractal
Here, single brackets denote summations over the nearest neighbors, and the double brackets – over the next-nearest neighbors.
H.2 $\text{NIPR}_{2}$ of states for the negative Vicsek lattice
In Fig. S23, we present the $\text{nIPR}_{2}$ values for all eigenstates in the energy spectrum of the fourth-generation NV fractal. The complete set of eigenstates and eigenvalues was obtained by exact diagonalization of the Hamiltonian in Eq. (18) with open boundary conditions. The parameters are set as $t=t^{\prime}=1$, $\bar{t}=\bar{t^{\prime}}=10^{-9}$, and $V=10^{4}$.
Figure S23: Energy spectrum of the single-particle tight-binding model on the fourth-generation negative Vicsek fractal. The color bar represents the normalized Inverse Participation Ratio ($\text{nIPR}_{2}$) for each eigenstate: blue and small markers signifies localized states (low $\text{nIPR}_{2}$), while red and big markers indicates ergodic/extended states (high $\text{nIPR}_{2}$). Although the large "clean" regions within the NV fractal structure might be expected to suppress the formation of NEMs, the fractal’s underlying geometric complexity still holds potential for the emergence of NEMs in certain spectral regions.

### v4

@@ E.2 Skewness
Skewness (Sk) quantifies the asymmetry of a probability distribution [84, 85]. For the wavefunction intensity, $|\psi_{E}|_{i}^{2}$, across lattice sites, Sk reveals deviations from a symmetric spatial spread. For instance, a perfectly uniform (ideal extended) state would have $\text{Sk}=0$. In contrast, localized states highly concentrated on a few sites with most other sites having near-zero intensity would exhibit large positive skewness. Skewness therefore serves as an additional qualitative indicator of spatial inhomogeneity, complementing the quantitative details of multifractal analysis. It is defined as:
@@
where $|\psi_{E}|_{i}^{2}$ is the wavefunction intensity at site $i$ for an eigenstate $\Phi_{E}$, $N$ is the total number of active sites for a given system generation $g$ (i.e. $N_{g}=N$), and the bracket denotes the spatial average.
@@
Fig. 20 presents the finite-size scaling (FSS) of the skewness, $\text{Sk}(N)$, for various characteristic states, obtained by plotting $\text{Sk}(N)$ against $1/\ln N$. The left panel of Fig. 20 establishes benchmark behaviors by displaying the FSS of skewness for the representative extended, localized, and multifractal states discussed in the MT. The extended ground state (GS, red circles) and the localized state (LOC, blue triangles) align with theoretical expectations: the former’s skewness values extrapolate towards $\text{Sk}_{\infty}\approx 0$ (consistent with its relatively uniform spatial spread), whereas the latter has a large, positive skewness that increases with system size $N$ (as for strong concentration on few sites). The representative non-ergodic multifractal state (NEM) also displays significant positive skewness for finite system sizes, trending towards a smaller yet distinctly positive $\text{Sk}_{\infty}$. Our analysis primarily uses this observed trend to differentiate the NEM’s behavior from both the near-zero skewness characteristic of extended states and the large skewness found for localized states. With this in mind, in the right panel of Fig. 20 we present the skewness FSS for the eigenstates at $E\simeq 0.000$, $E\simeq 1.000$, and $E\simeq 2.000$.
@@
Figure 20: Finite-size scaling of wavefunction intensity skewness $\text{Sk}(N)$ versus $1/\ln N$. (Left panel) Benchmark behaviors for representative extended (GS), multifractal (NEM), and localized (LOC) states. These states are extensively discussed in the Main Text. (Right panel) Skewness scaling for eigenstates at energies $E\simeq 0.000$, $E\simeq 1.000$, and $E\simeq 2.000$. Data points correspond to different system sizes and trends indicate the behavior towards the thermodynamic limit.
@@
The state at $E\simeq 0.000$ (green diamonds) is notable as it converges to a distinct, negative skewness, $\text{Sk}_{\infty}\simeq-1.055$ (determined from individual FSS fits not shown in this figure). Negative skewness typically implies that the distribution has a long tail towards lower intensity values: the majority of participating sites have higher-than-average intensities, with a sharp drop-off at the very low-intensity end. The amplitude profile of this state (Fig. 21a) visually corroborates this interpretation of its asymmetry; this observation, when considered alongside the mass exponent characterization, supports its classification as weakly ergodic.
@@
The eigenstate at $E\simeq 2.000$ (cyan hexagons) is also classified as weakly ergodic. Although its skewness is relatively stable and extrapolates to a small but positive $\text{Sk}_{\infty}$, its amplitude profile (Fig. 21c) reveals a broad, non-uniform distribution of intensity across many sites, with sharp probability decays occurring only at the edges of the larger groups of voids of the SG. These features, combined with the state’s high mass exponent value, point towards its classification also as weakly ergodic.
@@
Figure 21: Intensity profiles of $|\psi|^{2}$ corresponding states at energies (from left to right): $E\simeq 0.000$, $E\simeq 1.000$, and $E\simeq 2.000$. The states for $E\simeq 0.000$ and $E\simeq 2.000$ are weakly ergodic. The state for $E\simeq 1.000$ is a case of clustered-patterned multifractality. The color intensity, indicated by the colorbar, represents $|\psi|^{2}$ on a logarithmic scale.
@@
Finally, the state at $E\simeq 1.000$ (purple triangles) exhibits a clear positive extrapolated skewness. From the FSS, its $\text{Sk}(N)$ values drift towards a moderate positive value for larger system sizes. This trend, while distinct from those of the weakly ergodic states at $E\simeq 0.000$ and $E\simeq 2.000$, instead appears more similar to the benchmark NEM. Its probability distribution, while not uniform, is situated on a structured subset of sites with significant intensity fluctuations, leading to its positive skewness from extended regions of low probability. This aligns with its intensity profile (Fig. 21b), which reveals clusters of varying intensities around the g3g-3 sub-gasket blocks, supporting its characterization as a form of ’clustered-patterned multifractality’.
Appendix F Robustness
F.1 Local Robustness of Multifractality
To test whether the revealed NEMs are isolated states—requiring infinite precision to detect—we analyze the stability of the two-point correlator $\mathcal{F}(\omega,E)$ against small variations in energy. The results, shown in Fig. 22 for two spectral regions, reveal that while the overall power-law behavior is preserved across the energy neighborhood, quantitative details may vary between nearby states. This observation indicates that NEMs are robust spectral features forming a mini-band [13], rather than isolated occurrences.
Figure 22: Stability of the two-point correlator $\mathcal{F}(\omega,E)$ for states within a finite energy window. Analysis for generation $g=6$ states in the vicinity of the representative NEM at $E\simeq 1.509$ (left) and around $E\simeq-1.297$ (right). The persistence of the general power-law decay points to a robust multifractal character, though quantitative differences in plateau heights and decay exponents are apparent. The dashed line indicates the Gaussian Orthogonal Ensemble (GOE) value for ergodic states.
F.2 Robustness Against Randomly Displaced Defects
@@
Fig. 23 illustrates the evolution of the degeneracy for the special multifractal state at $E\simeq 1.000$ as a function of the number of randomly introduced defects. Contrary to the common expectation that adding defects would disrupt the fractal’s self-similar structure and lift the degeneracy of this energy level, an opposite trend is observed. The introduction of random defects may enhance certain quantum interference patterns, leading to an increased degeneracy. The changes to its spatial intensity profile are displayed in Fig. 24, which compares representative configurations before and after the addition of these random defects.
@@
Figure 23: Change in the average degeneracy $\langle n(E)\rangle$ of the multifractal state at $E\simeq 1.000$ in a $g=6$ SG as additional random defects are introduced. The plotted values are averages over $p=10$ distinct disorder realizations.
@@
Figure 24: Intensity profiles of the multifractal state at energy $E\simeq 1.000$ on a $g=6$ SG. Left panel: The state in the pristine system, without additional random defects (ARDs). Center and right panels: Two distinct realizations of the state when 150 ARDs are introduced. While some of the clustered features may persist, the introduction of ARDs can generate new quantum interference patterns, and the intensity is redistributed primarily towards the lower hierarchical levels of the SG structure.
F.3 Next-Nearest-Neighbor Hopping
@@
Figure 25: The frequency correlator $\mathcal{F}(\omega,E)$ for representative non-ergodic multifractal (NEM) states, illustrating the impact of varying next-nearest-neighbor (NNN) hopping strengths $t^{\prime}$. The plots reveal how characteristics such as plateau width and slope behavior respond to changes in $t^{\prime}$. Lines serve as guides to the eye.
Fig. 25 presents the behavior of $\mathcal{F}(\omega,E)$ for NEM states, specifically at energies $E\simeq 1.509$ and $E\simeq-1.287$, computed on a $g=6$ SG. The introduction of a small NNN hopping strength, up to $t^{\prime}=10^{-2}$, does not significantly modify the key qualitative features of the correlator, such as its plateau height and characteristic power-law decay exponent.
@@
Appendix G Generalization to Other Fractal Geometries and Interactions
An intriguing question arising from our findings is the extent to which the emergence of NEMs is specific to the SG, or if it represents a more general phenomenon. From our results, we believe NEMs can manifest in diverse fractal structures, though their characteristics would likely depend on intrinsic geometric properties such as the ramification number—a factor known to be significant in related studies on general graphs and fractals [54, 39].
In this spirit, investigating fractals embedded in three-dimensional space may offer a concrete pathway for observing NEM signatures [14]. However, performing comprehensive numerical analyses within three-dimensional configurations presents significant computational challenges, currently limiting extensive quantitative verification.
@@ Appendix G Generalization to Other Fractal Geometries and Interactions
G.1 Negative Vicsek Fractal
Conceptually, the negative Vicsek (NV) fractal is easier to fabricate than the SG, as it requires significantly fewer defects to be placed in a controllable manner. The starting point is a uniform square lattice $\mathcal{S}$ of $N$ sites. Then, unlike in the SG case, we retain sites corresponding to voids of the fractal and remove all other sites, see Figs. 26 and  27. We denote the set of indices of the removed sites as $\mathcal{B}\subset\mathcal{S}$.
Figure 26: Transformation of a minimal square lattice into the first-generation negative Vicsek fractal. Next-nearest-neighbor hopping is required to keep the lattice connected.
Figure 27: Transformation of a square lattice into the fourth-generation negative Vicsek fractal. The final structure emerges from successive iterations of the elementary transformation, where lattice sites in subset $\mathcal{B}$ are progressively replaced by defects. The resulting fractal is characterized by patching large regions of intact lattice without defects and regions with intricate defect patterns.
As in the SG case, each defect at position $r_{i}$ is modeled by a delta-peaked potential, $V_{i}=V\,\delta(r-r_{i})$, with a strength $V$ much larger than all other energy scales in the system. However, in the NV model, we also account for next-nearest-neighbor interactions, denoted by $t^{\prime}$. These additional couplings are fundamental, as they maintain connectivity between otherwise isolated regions of the fractal. Defect couplings are weak. The corresponding Hamiltonian reads:
$H=\sum\limits_{\begin{subarray}{c}\langle i,j\rangle\\ i,j\notin\mathcal{B}\end{subarray}}t\,c^{\dagger}_{i}c_{j}+\sum\limits_{\begin{subarray}{c}\langle\langle i,k\rangle\rangle\\ i,k\notin\mathcal{B}\end{subarray}}t^{\prime}\,c^{\dagger}_{i}c_{k}+\sum_{l\in\mathcal{B}}V\,c^{\dagger}_{l}c_{l}+\sum_{\begin{subarray}{c}\langle j,l\rangle\\ l\in\mathcal{B}\end{subarray}}\overline{t}\,c^{\dagger}_{j}c_{l}+\sum_{\begin{subarray}{c}\langle\langle k,l\rangle\rangle\\ l\in\mathcal{B}\end{subarray}}\overline{t^{\prime}}\,c^{\dagger}_{k}c_{l}+\text{h.c.}$ (19)
Here, single brackets denote summations over the nearest neighbors, and the double brackets over the next-nearest neighbors.
G.2 $\text{NPR}_{2}$ of States for the Negative Vicsek Lattice
@@
Figure 28: Energy spectrum of the single-particle tight-binding model on the fourth-generation negative Vicsek fractal. The colorbar represents the normalized Participation Ratio ($\text{nPR}_{2}$) for each eigenstate: small blue markers signify localized states (low $\text{nPR}_{2}$), while large red markers indicate ergodic/extended states (high $\text{nPR}_{2}$). Although the large "clean" regions within the NV fractal structure might be expected to suppress the formation of NEMs, the fractal’s underlying geometric complexity still holds potential for the emergence of NEMs in certain spectral regions.